OpenStudy (anonymous):

Suppose that the functions and are defined as follows q (x) = 2x r (x) = 2x^2 find the following (q o r ) (-1) = (r o q ) (-1) =

OpenStudy (anonymous):

(q o r)(x) = q(r(x)) = q(2x^2) = 2(2x^2) = 4x^2. Thus (q o r)(-1) = 4(-1)^2 = 4. (r o q)(x) = r(q(x)) = r(2x) = 2(2x)^2 = 2(4x^2) = 8x^2. Thus (r o q)(-1) = 8(-1)^2 = 8.

OpenStudy (anonymous):

didnt get it ?